OpenStudy (anonymous):

what is the simplified form of b/2/4/2b

OpenStudy (anonymous):

do you mean this? \[\frac{\frac{b}{2}}{\frac{4}{2b}}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if you're dividing by a fraction its the same as multiplying the recipocal example \[\frac{\frac{a}{b}}{\frac{c}{d}}=\frac{a}{b}*\frac{d}{c}=\frac{ad}{bc}\]

OpenStudy (anonymous):

\[b/2/4/2b=(((b/2)/4)/2)b=\frac{b^2}{16} \]